OpenStudy (anonymous):

when a force P act on a cart of mass 20kg. it moves with a uniform velocity on a horizontal floor. when a force 1.2p act on the cart ,it moves with an acceleration 0.05m/s^2. find the value of P?

OpenStudy (badhi):

Since when either P or 1.2P of force is applied, the object moves. Therefore the frictional force applied is constant (\(F_k\)) Apply \(F=ma\) to the two situations, \[P-F_k=m(0)=0 \\ 1.2P-F_k=ma=(20)(.05)\] FInd P